# Cosmic Ray Superdiffusion and Mirror Diffusion in Partially Ionized and Turbulent Medium

### v1

@@ 4.1.3 Magnetic field distribution and magnetic field spectrum
Fig. 4 presents 2D slices of the total magnetic field strength and corresponding magnetic energy spectra. In the strongly coupled regime, magnetic field fluctuations exhibit filamentary structures under transonic conditions, closely resembling the velocity structures shown in Fig. 2. The magnetic energy spectrum in this case generally follows a Kolmogorov scaling with a slope of approximately -5/3. Under supersonic conditions, numerous discontinuities associated with shocks are evident, resulting in a steeper magnetic energy spectrum characterized by a slope close to -2. Small-scale magnetic field structures become significantly less prominent in the intermediate coupled and decoupled regimes. Correspondingly, the magnetic energy spectra steepen beyond a slope of -2 for both supersonic and transonic conditions, indicating that magnetic energy becomes increasingly concentrated at larger scales. This behavior naturally arises due to substantial damping caused by neutral-ion collisions, effectively suppressing magnetic fluctuations.
4.2 Damping of magnetic field fluctuations increases CRs’ mean free path
@@
Damping caused by neutral-ion collisions effectively suppresses magnetic fluctuations at smaller scales, implying reduced pitch-angle scattering of CRs and consequently a larger mean free path. To quantify the mean free path of CR particles, we adopt the methodology described in Xu & Yan (2013). Specifically, we inject 1000 test particles at random initial positions, assigning initial cosine pitch angles ($\mu$) around 1, 0.25, and 0.1, respectively. The particle Larmor radius is set to $r_{L}=v/\Omega$ and is set to 10 cells, where $\Omega$ is the frequency of a particle’s gyromotion. We trace the change of pitch angle until it achieves 90 degrees. The corresponding averaged distance traveled along the local magnetic field is taken as the parallel mean free path.
@@
In Fig. 5, we present the mean free path $\langle\lambda_{\parallel}\rangle$ as functions of the dimensionless coupling strength $(\gamma_{\rm d}\rho_{i})/(v_{\rm inj}/L_{\rm inj})$, defined as the ratio of the neutral-ion collisional frequency to the eddy turnover rate. Under both supersonic and transonic conditions, an increase in $\langle\lambda_{\parallel}\rangle$ corresponds to weaker coupling, consistent with the suppression of small-scale magnetic fluctuations. When $\mu\approx 1$, the parallel motion of CRs along magnetic field lines is primarily governed by the resonant pitch-angle scattering and undergoes scattering diffusion. We trace the change of pitch angle until it achieves 90 degrees. The corresponding averaged distance traveled along the local magnetic field is taken as the parallel mean free path. However, at smaller pitch angles, mirror diffusion is dominant, and $\langle\lambda_{\parallel}\rangle$ is determined by the characteristic size of magnetic mirrors with which mirroring particles predominantly interact (Lazarian & Xu, 2021). As mirroring CRs repeatedly reverse their direction of motion along magnetic field lines, mirror diffusion strongly enhances CR confinement, resulting in a reduced mean free path at small $\mu$ (Zhang & Xu, 2023).
@@ 4.3 CRs’ perpendicular superdiffusion
To investigate perpendicular superdiffusion, we simultaneously inject 50 beams of test particles randomly distributed within the simulation cube, with each beam containing 20 particles. The initial spatial separation between particles within each beam is set to 10 grid cells. Consistent with the previous setup, all particles have a Larmor radius of $r_{L}=10$ cells and an initial $\mu\approx 1$, $\mu\approx 0.25$, or $\mu\approx 0.1$, respectively. We trace particle trajectories along local magnetic field lines at each time step and interpolate these trajectories to measure the perpendicular separation $\delta\widetilde{z}$ between particle trajectories at identical times. The root-mean-square (rms) value $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ is taken as the perpendicular displacement of the particles.
@@
Fig. 6 presents the correlation of $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ and the CRs’ gyro periods $t\Omega$ in transonic condition with initial $\mu\approx 1$. In these conditions, the parallel mean free path $\langle\lambda_{\parallel}\rangle$ exceeds the turbulence injection scale $L_{\rm inj}$ across all three coupling regimes. After surpassing the numerical dissipation scale, $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ exhibits a power-law growth with a slope of approximately. The results in supersonic conditions are similar, as studied by Hu et al. (2022). The growth of $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ is slower in the decoupling regime is slower due to the damping of small-scale magnetic fluctuations, which significantly reduces particle scattering.
@@
We repeat this analysis for particles with initial $\mu\approx 0.25$ and present the results in Fig. 6. In this scenario, mirror diffusion and scattering effects become important, reducing $\langle\lambda_{\parallel}\rangle$. In the decoupled regime, where $\langle\lambda_{\parallel}\rangle>L_{\rm inj}$, the perpendicular displacement $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ again follows a power-law scaling with a slope of approximately 3/2. However, in the intermediately and strongly coupled regimes, where damping of MHD turbulence is less significant, the power-law slope gradually deviates from 3/2, approaching a slope closer to 3/4. The power-law slope does not perfectly align with 3/4 because trajectories are averaged over 1000 particles. Since $\langle\lambda_{\parallel}\rangle$ is not significantly smaller than $L_{\rm inj}$, some particles with larger mean free paths still contribute to a slope closer to 3/2. Although the perpendicular displacement increases more slowly in these cases, the CR perpendicular motion remains superdiffusive. Additionally, $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ exhibits oscillations before entering the superdiffusive regime. These oscillations, absent in the $\mu\approx 1$ case, as shown in Fig. 6, could result from large magnetic mirrors, where CRs require more than one bounce for escaping (see Fig. 1). In the decoupled regime, small-scale magnetic fluctuations are damped, causing CRs to reflect predominantly from large-scale magnetic mirrors, thus producing more pronounced oscillations in $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$. The effects of interaction with mirrors of different sizes require further studies. Nevertheless, when the initial $\mu\approx 0.1$, the CRs’ $\langle\lambda_{\parallel}\rangle$ is smaller than $L_{\rm inj}$ in the three coupling/decoupling regimes. The power-law slope of $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ gradually deviates from 3/2, but is closer to 3/4. The oscillation induced by mirrors is still apparent.
@@ 5 Conclusion
In this work, we have examined the effects of partially ionized, turbulent environments on CR transport using high-resolution, 3D two-fluid simulations. By explicitly accounting for ion-neutral interactions, we have shown that neutral-ion collisions significantly alter turbulence dynamics, leading to important consequences for CR diffusion. Our main findings can be summarized as follows:
@@
    Neutral-ion decoupling and turbulence damping: As neutral-ion coupling weakens, we observe substantial damping of velocity, density, and magnetic field fluctuations at small scales, resulting in steeper turbulence spectra than classical Kolmogorov or Burgers scalings. Under supersonic conditions, shocks partially counteract this damping, sustaining density fluctuations and ionization enhancements, thereby influencing the coupling strength and turbulence structure.
@@
    Impact on CR parallel propagation: The damping of magnetic field fluctuations reduces pitch-angle scattering, substantially increasing the CR mean free path along the magnetic field direction. Importantly, mirror diffusion becomes increasingly significant at larger pitch angles, decreasing the parallel transport and increasing CRs confinement.
@@
    Perpendicular CR superdiffusion: We identify two distinct superdiffusive regimes of CR transport perpendicular to the mean magnetic field. In the strongly coupled regime, frequent scattering and mirror diffusion limit the parallel mean free path, and CR perpendicular transport follows a superdiffusive scaling of $\propto t^{3/4}$. Conversely, in the intermediately coupled and fully decoupled regimes, the suppression of small-scale magnetic fluctuations reduces scattering, leading to transport dominated by magnetic field line wandering, characterized by a scaling of $\propto t^{3/2}$.
Y.H. thanks Anatoly Spitkovsky for the helpful discussion. Y.H. acknowledges the support for this work provided by NASA through the NASA Hubble Fellowship grant No. HST-HF2-51557.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. A.L. acknowledges the support of NSF grants AST 2307840. This work used SDSC Expanse CPU and NCSA Delta CPU through allocations PHY230032, PHY230033, PHY230091, PHY230105, PHY230178, and PHY240183, from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.

### v2

@@ 4.1.3 Magnetic field distribution and magnetic field spectrum
Fig. 4 presents 2D slices of the total magnetic field strength and corresponding magnetic energy spectra. In the strongly coupled regime, magnetic field fluctuations exhibit filamentary structures under transonic conditions, closely resembling the velocity structures shown in Fig. 2. The magnetic energy spectrum in this case generally follows a Kolmogorov scaling with a slope of approximately -5/3. Under supersonic conditions, numerous discontinuities associated with shocks are evident, resulting in a steeper magnetic energy spectrum characterized by a slope close to -2. Small-scale magnetic field structures become significantly less prominent in the intermediate coupled and decoupled regimes. Correspondingly, the magnetic energy spectra steepen beyond a slope of -2 for both supersonic and transonic conditions, indicating that magnetic energy becomes increasingly concentrated at larger scales. This behavior naturally arises due to substantial damping caused by neutral-ion collisions, effectively suppressing magnetic fluctuations associated with the turbulent cascade.
4.2 Examine the damping effect on CR’s parallel mean free path
@@
Damping caused by neutral-ion collisions effectively suppresses magnetic fluctuations at smaller scales, implying reduced pitch-angle scattering of CRs and consequently a larger mean free path. To quantify the mean free path of CR particles, we adopt the methodology described in Xu & Yan (2013). Specifically, we inject 1000 test particles at random initial positions, assigning initial pitch angle cosine ($\mu$) around 1, 0.25, and 0.1, respectively. The particle Larmor radius $r_{L}=v/\Omega$ is set to 10 cells, where $\Omega$ is the particle’s gyrofrequency. We trace the change of pitch angle until it has changed by 90 degrees. The corresponding averaged distance traveled along the local magnetic field is taken as the parallel mean free path.
In Fig. 5, we present the mean free path $\langle\lambda_{\parallel}\rangle$ average over all particles as functions of the dimensionless coupling strength $(\gamma_{\rm d}\rho_{i})/(v_{\rm inj}/L_{\rm inj})$, defined as the ratio of the neutral-ion collisional frequency to the eddy turnover rate at injection scale.
@@
In the strong-coupling regime, for CRs with initially large $\mu$, i.e., small pitch angles, the parallel motion of CRs along magnetic field lines is initially governed by the resonant pitch-angle scattering and undergoes scattering diffusion until the pitch angle becomes sufficiently large, and they enter the mirror diffusion regime. By contrast, for CRs with initially small $\mu$, i.e., large pitch angles, mirror diffusion is dominant, and $\langle\lambda_{\parallel}\rangle$ is determined by the characteristic size of magnetic mirrors with which mirroring particles predominantly interact (Lazarian & Xu, 2021). As mirroring of CRs effectively reverses their direction of motion along magnetic field lines, mirror diffusion strongly enhances CR confinement, resulting in a reduced mean free path at initially small $\mu$ (Zhang & Xu, 2023).
Nevertheless, under both supersonic and transonic conditions, $\langle\lambda_{\parallel}\rangle$ increases with decreasing coupling strength, consistent with stronger damping of MHD turbulence. Damping reduces pitch-angle scattering efficiency by suppressing magnetic fluctuations at gyroradius scales. However, the resulting increase in $\langle\lambda_{\parallel}\rangle$ is not significant—less than an order of magnitude—indicating that even with substantial damping and suppression of small-scale fluctuations, CRs at large pitch angles remain well confined by non-resonant mirror diffusion. In this regime, $\langle\lambda_{\parallel}\rangle$ can remain smaller than $L_{\rm inj}$ and the cloud size, ensuring that CRs are dynamically coupled to the cold ISM.
@@ 4.3 CRs’ perpendicular superdiffusion
To investigate perpendicular superdiffusion, we follow the approach in Hu et al. (2022), simultaneously injecting 50 beams of test particles randomly distributed within the simulation cube, with each beam containing 20 particles. The initial spatial separation between particles within each beam is set to 10 grid cells. Consistent with the previous setup, all particles have a Larmor radius of $r_{L}=10$ cells and an initial $\mu\approx 1$, $\mu\approx 0.25$, or $\mu\approx 0.1$, respectively. We trace particle trajectories along local magnetic field lines at each time step and interpolate these trajectories to measure the separation perpendicular to the mean magnetic field $\delta\widetilde{z}$ between particle trajectories at identical times. The root-mean-square (rms) value $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ is taken as the perpendicular displacement of the particles.
@@
Fig. 6 presents the time evolution of $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ in transonic condition with initial $\mu\approx 1$. In these conditions, the parallel mean free path $\langle\lambda_{\parallel}\rangle$ exceeds the turbulence injection scale $L_{\rm inj}$ across all three coupling regimes. After surpassing the numerical dissipation scale, $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ exhibits a power-law growth with a slope of approximately 3/2. When $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ is larger than $L_{\rm inj}$, the growth slows down, showing a slope of approximately 1/2. The results in supersonic conditions are similar, as studied by Hu et al. (2022). In the decoupled regime, however, $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ grows more slowly before entering the perpendicular superdiffusion regime, requiring approximately 1000 gyroperiods compared to only 10 gyroperiods in the strongly coupled case. This delay arises because the initial perpendicular separation of CRs is smaller than the maximum parallel decoupling scale, and the damping of small-scale magnetic fluctuations modifies the superdiffusive behavior of magnetic field lines. The delay also makes it difficult to see the change of slope beyond $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}>L_{\rm inj}$. As the required computational time increases exponentially by orders of magnitude, we stop the simulation at $10^{4}~~t\Omega$, while longer time may be investigated in future work.
@@
We repeat this analysis for particles with initial $\mu\approx 0.25$ and present the results in Fig. 6. In this scenario, mirror diffusion becomes more important, reducing $\langle\lambda_{\parallel}\rangle$. In the decoupled regime, where $\langle\lambda_{\parallel}\rangle>L_{\rm inj}$, the perpendicular separation $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ again follows a power-law scaling with a slope of approximately 3/2. However, in the intermediate and strongly coupled regimes, where damping of MHD turbulence is less significant, the power-law slope gradually deviates from 3/2, approaching a slope closer to 3/4. The power-law slope does not perfectly align with 3/4 because trajectories are averaged over 1000 particles. Since $\langle\lambda_{\parallel}\rangle$ is not significantly smaller than $L_{\rm inj}$, some particles with larger mean free paths still contribute to a slope closer to 3/2. Although the perpendicular separation increases more slowly in these cases, the CR perpendicular motion remains superdiffusive. Additionally, $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ exhibits oscillations before entering the superdiffusive regime. These oscillations, absent in the $\mu\approx 1$ case, as shown in Fig. 6, could result from the perpendicular displacement associated with the CRs’ initial gyro-orbit. In the decoupled regime, small-scale magnetic fluctuations are damped, causing CRs to move along the magnetic field line ballistically, thus producing more pronounced oscillations in $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$. Nevertheless, when the initial $\mu\approx 0.1$, the CRs’ $\langle\lambda_{\parallel}\rangle$ is smaller than $L_{\rm inj}$ in the three coupling/decoupling regimes, and the oscillation possibly induced by mirrors is still apparent. The power-law slope of $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ gradually deviates from 3/2, but is closer to 3/4. This finding is similar to that in Zhang & Xu (2023), who measured the perpendicular superdiffusion of CRs with initially large pitch angles in a single-fluid MHD turbulence simulation. It shows that large-pitch-angle CRs are well confined in turbulence, irrespective of the coupling state of neutrals to ions.
@@ 5 Conclusion
In this work, we have examined the effects of partially ionized, turbulent environments on CR transport using high-resolution, 3D two-fluid simulations. By explicitly accounting for ion-neutral interactions, we have shown that despite the ion-neutral collisional damping on small-scale turbulent magnetic fluctuations, the damping effect on CR transport is not significant. Our main findings can be summarized as follows:
@@
    Neutral-ion decoupling and turbulence damping: As neutral-ion coupling weakens, we observe substantial damping of velocity, density, and magnetic field fluctuations at small scales, resulting in steeper turbulence kinetic spectra than classical Kolmogorov or Burgers scalings. Under supersonic conditions, large-scale shocks do not suffer damping, sustaining the small-scale density enhancements due to the shock compression, thereby influencing turbulence structure and coupling strength, which is defined as the ratio of neutral-ion collisional frequency and eddy turnover time.
@@
    Impact on CR parallel propagation: In the strongly coupled regime, frequent scattering and mirror diffusion can limit the parallel mean free path. Mirror diffusion becomes increasingly significant at larger pitch angles, decreasing the parallel mean free path and improving CR confinement. In the intermediately coupled and fully decoupled regimes, although the damping of small-scale magnetic fluctuations reduces resonant pitch angle scattering, the change of parallel mean free path of CRs is less than an order of magnitude, due to their non-resonant interactions with turbulent mirrors. The results imply that large-pitch-angle CRs can be well confined in the cold ISM, such as molecular clouds.
@@
    Perpendicular CR superdiffusion: We identify two distinct superdiffusive regimes of CR transport perpendicular to the mean magnetic field: a diffusive regime, where the CRs’ parallel mean free path is smaller than the turbulence injection scale, and a ballistic-like regime, where the parallel mean free path is larger than the injection scale. In the diffusive regime, CRs’ perpendicular transport follows a superdiffusive scaling of $\propto t^{3/4}$. Conversely, in the ballistic-like regime, CRs’ perpendicular diffusion is characterized by a scaling of $\propto t^{3/2}$.
Y.H. thanks Anatoly Spitkovsky for the helpful discussion. Y.H. acknowledges the support for this work provided by NASA through the NASA Hubble Fellowship grant No. HST-HF2-51557.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. S.X. acknowledges the support from the NASA ATP award 80NSSC24K0896. A.L. acknowledges the support of NSF grants AST 2307840. This work used SDSC Expanse CPU and NCSA Delta CPU through allocations PHY230032, PHY230033, PHY230091, PHY230105, PHY230178, and PHY240183, from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.